# Robust First and Second-Order Differentiation for Regularized Optimal Transport

### v1

@@ Appendix A Proofs for spectral analysis
Let $\Pi,\Pi^{*}$ be two positive coupling matrices with $A=\Pi-\Pi^{*}$ satisfying
@@
Then, the eigenvalues of their H-matrices are close:
$|\lambda_{k}(H(\Pi))-\lambda_{k}(H(\Pi^{*}))|\leq\delta_{1}+\delta_{2},\quad 1\leq k\leq N+M.$
@@
$H(\Pi)-H(\Pi^{*})=\left[\begin{array}[]{ccccccccc}\mathrm{diag}(\sum_{j=1}^{N}A_{i,j})&A\\ A^{\top}&\mathrm{diag}(\sum_{i=1}^{N}A_{i,j})\end{array}\right]=:E.$ (64)
By Weyl’s inequality, we have $|\lambda_{k}(H^{*})-\lambda_{k}(H)|\leq\|E\|_{op}$. Thus, it suffices to estimate $\|E\|_{op}$. Note that first that using $|A_{i,j}|\leq 1$ and (61), we have $\sum_{i}|\sum_{j}A_{i,j}u_{i}|^{2}=\sum_{i}|\sum_{j}A_{i,j}|^{2}u_{i}^{2}\leq\sum_{i}|\sum_{j}A_{i,j}|u_{i}^{2}\leq\delta_{1}^{2}\|u\|^{2},$ and similarly, $\sum_{j}|\sum_{i}A_{i,j}v_{j}|^{2}\leq\delta_{1}^{2}\|v\|^{2}$; also, $\|Av\|^{2}=\sum_{i=1}|\sum_{j}A_{i,j}v_{j}|^{2}\leq\sum_{i=1}[\sum_{j}A_{i,j}^{2}\sum_{j}|v_{j}|^{2}]\leq\delta_{2}^{2}\|v\|^{2},$ and similarly, $\|A^{\top}u\|^{2}\leq\delta_{2}^{2}\|u\|^{2}$.
Using these four bounds, we have $\|E\|_{op}^{2}=\sup_{\boldsymbol{u}\in\mathbb{R}^{M},\boldsymbol{v}\in\mathbb{R}^{N},\|\boldsymbol{u}\|^{2}+\|\boldsymbol{v}\|^{2}=1}\|E\left[\begin{array}[]{ccccccccc}\boldsymbol{u}\\ \boldsymbol{v}\end{array}\right]\|^{2}$, then $\|E\|_{op}^{2}=\sum_{i}|\sum_{j}A_{i,j}u_{i}|^{2}+\|Av\|^{2}+\sum_{j}|\sum_{i}A_{i,j}v_{j}|^{2}+\|A^{\top}u\|^{2}\leq\delta_{1}^{2}+\delta_{2}^{2}\leq\delta_{1}+\delta_{2}$. Combining with Weyl’s inequality, we conclude the proof. ∎
@@
The bound for the eigenvalues in (56) follows from (55) and Lemma 13. To prove the bounds for the condition number, recall that in the proof of Theorem 9, we have shown that $\lambda_{1}:=\lambda_{1}(H(\Pi^{*}))=\frac{1}{N}+\frac{1}{M}$ and $\lambda_{N+M-1}:=\lambda_{N+M-1}(H(\Pi^{*}))=\frac{1}{2}\lambda_{1}\left[1-\sqrt{1-4\Delta}\right]\in[\lambda_{1}\Delta,2\lambda_{1}\Delta]$ with $\Delta=\lambda_{1}^{-2}(\sigma_{1}^{2}-\sigma_{2}^{2})$. For $\delta+\delta_{2}=t\lambda_{1}\Delta$ with $t\in[0,1)$, Eq.(56) implies that
$(1-t)\lambda_{1}\Delta\leq\lambda_{N+M-1}-\delta-\delta_{2}$ $\leq\lambda_{N+M-1}(H(\Pi))\leq\lambda_{N+M-1}+\delta+\delta_{2}$
@@
$\lambda_{1}(1-t\Delta)\leq\lambda_{1}-(\delta+\delta_{2})$ $\leq\lambda_{1}(H(\Pi))\leq\lambda_{1}+\delta+\delta_{2}\leq\lambda_{1}(1+t\Delta).$
@@
$\frac{1+t\Delta}{(2+t)\Delta}\leq\frac{\lambda_{1}-(\delta+\delta_{2})}{\lambda_{N+M-1}+(\delta+\delta_{2})}\leq\kappa(H(\Pi))\leq\frac{\lambda_{1}+(\delta+\delta_{2})}{\lambda_{N+M-1}-(\delta+\delta_{2})}\leq\frac{1+t\Delta}{(1-t)\Delta}.$
Proof of Example 12.
Part (a): Note that the cost matrix $C$ is $C_{ij}=\|{y}_{i}-{y}_{j}\|_{2}^{2}=4\sin^{2}\left(x_{|j-i|}/2\right)$ and the Gibbs kernel $K_{ij}=\exp\left(-\frac{4\sin^{2}\left(|i-j|\pi/N\right)}{\epsilon}\right)$ is circulant, whose rows and columns sum is $\lambda_{1}(K)=\sum_{j=0}^{N-1}\exp\left(-\frac{4\sin^{2}\left(j\pi/N\right)}{\epsilon}\right)$. Then the matrix $\Pi^{*}=\frac{K}{\lambda_{1}(K)N}$ satisfies the uniform marginal constraints on $\Pi^{*}$, hence, it is the optimal coupling matrix due to uniqueness of the solution of the contraint optimization (4).
Part (b) follows directly from Corollary 10.
Part (c), We first compute the largest two eigenvalues of $K$. Recall that the matrix $K$ is symmetric and positive-definite, so its singular values are the same as its eigenvalues. Since $K$ is circulant, the first two eigenvalues of $K$ are (see e.g., [16]), we have
$\lambda_{1}(K)=\sum_{j=0}^{N-1}\exp\left(-\frac{4\sin^{2}(\frac{j\pi}{N})}{\epsilon}\right),\lambda_{2}(K)=\sum_{j=0}^{N-1}\exp\left(-\frac{4\sin^{2}(\frac{j\pi}{N})}{\epsilon}\right)\cos\left(\frac{2j\pi}{N}\right)$
Meanwhile, combining Part (a) and Part (b), we have $\lambda_{2N-1}(H)=\frac{1}{N}(1-\frac{\lambda_{2}(K)}{\lambda_{1}(K)})$. Thus, to study the limits, we first study the limit of $\lambda_{1}(K)$ and $\lambda_{2}(K)$.
@@
where $I_{1}(x)$ and $I_{2}(x)$ are the modified Bessel functions of first kind. Then, the limit of the second smallest eigenvalue of $H$ is
$\lim_{N\rightarrow+\infty}(N\cdot\lambda_{2N-1}(H))$ $=1-\lim_{N\rightarrow+\infty}\frac{\lambda_{2}(K)}{\lambda_{1}(K)}=1-\frac{I_{1}(2/\epsilon)}{I_{0}(2/\epsilon)}.$
@@
$\lim_{N\rightarrow+\infty}(N\cdot\lambda_{2N-1}(H))=1-\frac{\frac{(\epsilon/2)^{1/2}}{\sqrt{2\pi}}-\frac{3(\epsilon/2)^{3/2}}{8\sqrt{2\pi}}+O(\epsilon/2)^{5/2}}{\frac{(\epsilon/2)^{1/2}}{\sqrt{2\pi}}+\frac{(\epsilon/2)^{3/2}}{8\sqrt{2\pi}}+O(\epsilon/2)^{5/2}}=\frac{\epsilon}{4}+O(\epsilon^{2}).$
Then, we have $\lim_{\epsilon\rightarrow 0^{+}}\lim_{N\rightarrow+\infty}\frac{N\cdot\lambda_{2N-1}(H)}{\epsilon}=\frac{1}{4}$, and $\lim_{\epsilon\rightarrow 0^{+}}\lim_{N\rightarrow+\infty}\epsilon\frac{\lambda_{1}(H)}{\lambda_{2N-1}(H)}=8$, which gives (57).
To prove (58), note first that when $N$ is fixed and $\epsilon\rightarrow 0^{+}$, $\lambda_{1}(K)$ and $\lambda_{2}(K)$ are approximated by the three largest terms,
$\lambda_{1}(K)=1+2\exp(-\frac{4}{\epsilon}\sin^{2}(\frac{\pi}{N}))+O_{\epsilon},\lambda_{2}(K)=1+2\exp(-\frac{4}{\epsilon}\sin^{2}(\frac{\pi}{N}))\cos(\frac{2\pi}{N})+O_{\epsilon}$
@@
$\exp\left(\frac{4}{\epsilon}\sin^{2}(\frac{\pi}{N})\right)N\lambda_{2N-1}(H)=\frac{2(1-\cos(\frac{2\pi}{N}))+O_{\epsilon}}{1+2\exp\left(-\frac{4}{\epsilon}\sin^{2}(\frac{\pi}{N})\right)+O_{\epsilon}}.$
@@
$\lim_{N\rightarrow+\infty}\lim_{\epsilon\rightarrow 0^{+}}N^{3}\exp\left(\frac{4\sin^{2}(\pi/N)}{\epsilon}\right)\lambda_{2N-1}(H)=4\pi^{2}.$

### v2

@@ A.2 Proof of Proposition 12
Let $\bm{\Pi},\bm{\Pi}^{*}$ be two positive coupling matrices with $\mathbf{A}=\bm{\Pi}-\bm{\Pi}^{*}$ satisfying
@@
Then, the eigenvalues of their $\mathbf{H}$ matrices are close:
$|\lambda_{k}(\mathbf{H})-\lambda_{k}(\mathbf{H}^{*})|\leq\delta_{1}+\delta_{2},\quad 1\leq k\leq N+M.$
@@
$\mathbf{H}-\mathbf{H}^{*}=\left[\begin{array}[]{ccccccccc}\mathrm{diag}(\mathbf{A}\mathbb{1})&\mathbf{A}\\ \mathbf{A}^{\top}&\mathrm{diag}(\mathbf{A}^{\top}\mathbb{1})\end{array}\right]=:\mathbf{E}.$ (58)
By Weyl’s inequality, we have $|\lambda_{k}(\mathbf{H}^{*})-\lambda_{k}(\mathbf{H})|\leq\|\mathbf{E}\|_{op}$. Thus, it suffices to estimate $\|\mathbf{E}\|_{op}$. Note that first that using $|A_{ij}|\leq 1$ and (55), we have $\sum_{i}|\sum_{j}A_{ij}u_{i}|^{2}=\sum_{i}|\sum_{j}A_{ij}|^{2}u_{i}^{2}\leq\sum_{i}|\sum_{j}A_{ij}|u_{i}^{2}\leq\delta_{1}^{2}\|\mathbf{u}\|^{2},$ and similarly, $\sum_{j}|\sum_{i}A_{ij}v_{j}|^{2}\leq\delta_{1}^{2}\|\mathbf{v}\|^{2}$; also, $\|\mathbf{A}\mathbf{v}\|^{2}=\sum_{i=1}|\sum_{j}A_{ij}v_{j}|^{2}\leq\sum_{i=1}[\sum_{j}A_{ij}^{2}\sum_{j}|v_{j}|^{2}]\leq\delta_{2}^{2}\|\mathbf{v}\|^{2},$ and similarly, $\|\mathbf{A}^{\top}\mathbf{u}\|^{2}\leq\delta_{2}^{2}\|\mathbf{u}\|^{2}$.
Using these four bounds, we have $\|\mathbf{E}\|_{op}^{2}=\sup_{\mathbf{u}\in\mathbb{R}^{M},\mathbf{v}\in\mathbb{R}^{N},\|\mathbf{u}\|^{2}+\|\mathbf{v}\|^{2}=1}\|\mathbf{E}\left[\begin{array}[]{ccccccccc}\mathbf{u}\\ \mathbf{v}\end{array}\right]\|^{2}$, then $\|\mathbf{E}\|_{op}^{2}=\sum_{i}|\sum_{j}A_{ij}u_{i}|^{2}+\|\mathbf{Av}\|^{2}+\sum_{j}|\sum_{i}A_{ij}v_{j}|^{2}+\|\mathbf{A^{\top}u}\|^{2}\leq\delta_{1}^{2}+\delta_{2}^{2}\leq\delta_{1}+\delta_{2}$. Combining with Weyl’s inequality, we conclude the proof. ∎
@@
The bound for the eigenvalues in (50) follows from (49) and Lemma 14. To prove the bounds for the condition number, we have shown that $\lambda_{1}:=\lambda_{1}(\mathbf{H}^{*})=\frac{1}{N}+\frac{1}{M}$ and $\lambda_{N+M-1}:=\lambda_{N+M-1}(\mathbf{H}^{*})=\frac{1}{2}\lambda_{1}\left[1-\sqrt{1-4\Delta}\right]\in[\lambda_{1}\Delta,2\lambda_{1}\Delta]$ with $\Delta=\lambda_{1}^{-2}\rho(\bm{\Pi}^{\top}\bm{\Pi})$. For $\delta+\delta_{2}=t\lambda_{1}\Delta$ with $t\in[0,1)$, Eq.(50) implies that
$(1-t)\lambda_{1}\Delta\leq\lambda_{N+M-1}-\delta-\delta_{2}$ $\leq\lambda_{N+M-1}(\mathbf{H})\leq\lambda_{N+M-1}+\delta+\delta_{2}$
@@
$\lambda_{1}(1-t\Delta)\leq\lambda_{1}-(\delta+\delta_{2})$ $\leq\lambda_{1}(\mathbf{H})\leq\lambda_{1}+\delta+\delta_{2}\leq\lambda_{1}(1+t\Delta).$
@@
$\frac{1+t\Delta}{(2+t)\Delta}\leq\frac{\lambda_{1}-(\delta+\delta_{2})}{\lambda_{N+M-1}+(\delta+\delta_{2})}\leq\kappa(\mathbf{H})\leq\frac{\lambda_{1}+(\delta+\delta_{2})}{\lambda_{N+M-1}-(\delta+\delta_{2})}\leq\frac{1+t\Delta}{(1-t)\Delta}.$
Appendix B Equally spaced points on the unit circle
In this section, we will derive the precise asymptotic behavior of the spectral properties of $\mathbf{H}^{*}$ as a function of the regularization strength $\epsilon$ and the sample size $N$ in Example 4.8.
Proof.
Part (a): Note that the cost matrix is $C_{ij}=\|{y}_{i}-{y}_{j}\|_{2}^{2}=4\sin^{2}\left(x_{|j-i|}/2\right)$ and the Gibbs kernel $K_{ij}=\exp\left(-\frac{4\sin^{2}\left(|i-j|\pi/N\right)}{\epsilon}\right)$ is circulant, whose rows and columns sum is $\lambda_{1}(\mathbf{K})=\sum_{j=0}^{N-1}\exp\left(-\frac{4\sin^{2}\left(j\pi/N\right)}{\epsilon}\right)$. Then the matrix $\bm{\Pi}^{*}=\frac{\mathbf{K}}{\lambda_{1}(\mathbf{K})N}$ satisfies the uniform marginal constraints on $\bm{\Pi}^{*}$, hence, it is the optimal coupling matrix due to uniqueness of the solution of the constraint optimization.
Part (b) follows directly from Corollary 4.6.
Part (c), We first compute the largest two eigenvalues of $\mathbf{K}$. Recall that the matrix $\mathbf{K}$ is symmetric and positive-definite, so its singular values are the same as its eigenvalues. Since $\mathbf{K}$ is circulant, the first two eigenvalues of $\mathbf{K}$ are (see e.g., [18]), we have
$\lambda_{1}(\mathbf{K})=\sum_{j=0}^{N-1}\exp\left(-\frac{4\sin^{2}(\frac{j\pi}{N})}{\epsilon}\right),\lambda_{2}(\mathbf{K})=\sum_{j=0}^{N-1}\exp\left(-\frac{4\sin^{2}(\frac{j\pi}{N})}{\epsilon}\right)\cos\left(\frac{2j\pi}{N}\right)$
Meanwhile, combining Part (a) and Part (b), we have $\lambda_{2N-1}(\mathbf{H}^{*})=\frac{1}{N}(1-\frac{\lambda_{2}(\mathbf{K})}{\lambda_{1}(\mathbf{K})})$. Thus, to study the limits, we first study the limit of $\lambda_{1}(\mathbf{K})$ and $\lambda_{2}(\mathbf{K})$.
@@ Appendix B Equally spaced points on the unit circle
[…]
where $I_{1}(x)$ and $I_{2}(x)$ are the modified Bessel functions of first kind. Then, the limit of the second smallest eigenvalue of $\mathbf{H}^{*}$ is
$\lim_{N\rightarrow+\infty}(N\cdot\lambda_{2N-1}(\mathbf{H}^{*}))$ $=1-\lim_{N\rightarrow+\infty}\frac{\lambda_{2}(\mathbf{K})}{\lambda_{1}(\mathbf{K})}=1-\frac{I_{1}(2/\epsilon)}{I_{0}(2/\epsilon)}.$
@@
$\lim_{N\rightarrow+\infty}(N\cdot\lambda_{2N-1}(\mathbf{H}^{*}))=1-\frac{\frac{(\epsilon/2)^{1/2}}{\sqrt{2\pi}}-\frac{3(\epsilon/2)^{3/2}}{8\sqrt{2\pi}}+O(\epsilon/2)^{5/2}}{\frac{(\epsilon/2)^{1/2}}{\sqrt{2\pi}}+\frac{(\epsilon/2)^{3/2}}{8\sqrt{2\pi}}+O(\epsilon/2)^{5/2}}=\frac{\epsilon}{4}+O(\epsilon^{2}).$
Then, we have $\lim_{\epsilon\rightarrow 0^{+}}\lim_{N\rightarrow+\infty}\epsilon\frac{\lambda_{1}(\mathbf{H}^{*})}{\lambda_{2N-1}(\mathbf{H}^{*})}=8$, which gives (51).
To prove (52), note first that when $N$ is fixed and $\epsilon\rightarrow 0^{+}$, $\lambda_{1}(\mathbf{K})$ and $\lambda_{2}(\mathbf{K})$ are approximated by the three largest terms,
$\lambda_{1}(\mathbf{K})=1+2\exp(-\frac{4}{\epsilon}\sin^{2}(\frac{\pi}{N}))+O_{\epsilon},\lambda_{2}(\mathbf{K})=1+2\exp(-\frac{4}{\epsilon}\sin^{2}(\frac{\pi}{N}))\cos(\frac{2\pi}{N})+O_{\epsilon}$
@@
$\exp\left(\frac{4}{\epsilon}\sin^{2}(\frac{\pi}{N})\right)N\lambda_{2N-1}(\mathbf{H}^{*})=\frac{2(1-\cos(\frac{2\pi}{N}))+O_{\epsilon}}{1+2\exp\left(-\frac{4}{\epsilon}\sin^{2}(\frac{\pi}{N})\right)+O_{\epsilon}}.$
@@
$\lim_{N\rightarrow+\infty}\lim_{\epsilon\rightarrow 0^{+}}N^{3}\exp\left(\frac{4\sin^{2}(\pi/N)}{\epsilon}\right)\lambda_{2N-1}(\mathbf{H}^{*})=4\pi^{2}.$